# Linear constellations in primes with arithmetic restrictions

### v2

@@ Acknowledgements
We thank Pierre-Yves Bienvenu for helpful discussions. Most of this work was completed while M.T. was a visiting scholar at TU Graz in 2022–2024. The authors thank TU Graz for its hospitality. C.F. was supported by EPSRC grants EP/T01170X/1 and EP/T01170X/2. M.T. was supported by Czech Science Foundation (GAČR) grant 22-11563O.
@@ 2.1. Some notation
As already stated, for $N\in\mathbb{N}$ we let $[N]:=\{1,\ldots,N\}$. We think of number fields as subfields of $\mathbb{C}$. We denote the discriminant of a number field $K$ by $\Delta_{K}$, the monoid of non-zero ideals of its ring of integers $\mathscr{O}_{K}$ by $\mathscr{I}_{K}$, the maximal abelian subextionsion of $K$ over $\mathbb{Q}$ by $K^{\operatorname{ab}}$, and the finite part of the conductor of this abelian subfield by $\Phi_{K^{\operatorname{ab}}}$. We denote the absolute norm of $\mathfrak{a}\in\mathscr{I}_{K}$ by $\mathfrak{N}\mathfrak{a}$.
@@ 2.3. Polynomial subsequences of equidistributed nilsequences
We prove our main results using the machinery developed by Green, Tao and Ziegler, in particular the inverse theorem for Gowers norms [16]. This makes it necessary to deal with polynomial sequences in nilmanifolds. Green and Tao have established the quantitative theory of polynomial nilsequences in [15]. We use the definitions and terminology of this work freely and will refer to it frequently. In particular, we use Mal’cev bases as defined in [15, Definition 2.1] and the metrics induced by these Mal’cev bases as defined in [15, Definition 2.2]. Lipschitz constants are understood with respect to this metric.
In Section 4, we compile some facts regarding polynomial subsequences of equidistributed nilsequences, which are probably well known to experts. We have included full proofs, as we were not able to find the exact required statements in the literature.
@@ 2.4. W-trick
A central aspect in the strategy of Green, Tao and Ziegler is the $W$-trick, originating from [11]. It requires us to consider an increasing function $w(N)$ that grows to infinity with $N$ and satisfies $1\leq w(N)\leq\frac{1}{3}\log\log N$, a constant $\mathscr{D}\in\mathbb{N}$, and the function
@@
$\Lambda^{\prime}(n)$ :=$:=\begin{cases}\log p,&\text{ if }n=p\text{ prime},\\ 0,&\text{ otherwise,}\end{cases}$
$\Lambda_{K,C}^{\prime}(n)$ :=$:=\begin{cases}\log p,&\text{ if }n=p\text{ prime with $p$ unramified in $K$ and $[K/\mathbb{Q},p]=C$},\\ 0,&\text{ otherwise,}\end{cases}$
$\Lambda_{a}^{\prime}(n)$ :=$:=\begin{cases}\log p,&\text{ if }n=p\text{ prime with }\mathbb{F}_{p}^{\times}=\langle a\rangle,\\ 0,&\text{ otherwise.}\end{cases}$
@@
We will convince ourselves (Lemma 9.1) that the pseudorandom majorant constructed by Green and Tao in [13, Proposition 6.4] for $\Lambda_{b,W}$ is also sufficient for our $\Lambda_{K,C,b,W}$ and $\Lambda_{a,b,W}$. Standard arguments, including the generalised von Neumann Theorem [13, Proposition 7.1] (see Lemma 9.4), then reduce Theorem 1.4 to the following Gowers norm estimate. For the definition of $\lVert\cdot\rVert_{U^{u+1}[N]}$ we refer to [13, (B.13)].
@@
which was proved in [13, Theorem 7.2] with the choice of $\mathscr{D}=1$ and $w(N)=\log\log\log N$ but holds also in general (Lemma 9.2).
@@ 2.5. Von Mangoldt model and orthogonality with nilsequences
Green and Tao use the inverse theorem for Gowers norms [17] to deduce (2.3) from an orthogonality result of $\Lambda_{b,W}-1$ with nilsequences ([13, Proposition 10.2]). The proof of this result uses the identity
@@
Starting from this identity, Green and Tao reduce [13, Proposition 10.2] to a sieve estimate [13, (12.5)] and a strong orthogonality result for the Möbius function $\mu$ with nisequences, which they prove in [14, Theorem 1.1]. Unfortunately, an analog of (2.4) is not available for our functions $\Lambda_{K,C}$ and $\Lambda_{a}$.
Therefore, we proceed in a different way in our proofs of Proposition 2.1 and Proposition 2.2: in order to show orthogonality with nilsequences, we apply directly the strategy used for the Möbius function in [14]. This makes crucial use of the fact that $\mu(n)$ is equidistributed in arithmetic progressions, which is clearly not entirely true for our function $\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)$ due to progressions such as $n\equiv 0\bmod b$. To deal with this issue, we replace the constant average $\phi(W)\eta_{K,C}(b,W)$ by a suitably scaled version of the $W$-tricked classical von Mangoldt function $\Lambda_{b,W}(n)$ and show that this provides a good model for $\Lambda_{K,C,b,W}(n)$ in all arithmetic progressions. With this model, we show the following orthogonality result, which is analogous to [14, Theorem 1.1].
@@
$\ll_{K,m,d,A}Q^{O_{m,d,A}(1)}(1+\lVert F\rVert_{\operatorname{Lip}})(\log N)^{-A}.$
@@
$\Lambda_{K}(\mathfrak{a})=\begin{cases}\log\mathfrak{N}\mathfrak{p}&\text{ if }\mathfrak{a}=\mathfrak{p}^{e}\text{ with $\mathfrak{p}$ prime ideal and $e\in\mathbb{N}$},\\ 0&\text{ otherwise.}\end{cases}.$ (2.5)
@@
We prove Theorem 2.4 in §5 and deduce Theorem 2.3 from it in §6. Let us just mention in passing that it would also be possible to deal with $\Lambda_{K,C}$ in terms of the von Mangoldt function $\Lambda_{L}$ for ideals in certain subfields $L$ of $K$ (see Lemma 6.2). For these ideal von Mangoldt functions, an analogue of (2.4) exists and our techniques from §5 would be capable of also proving an analogue of [14, Theorem 1.1] for the Möbius function $\mu_{L}$ of ideals of $L$. However, it seems quite difficult to prove the required analogue of the sieve estimate [13, (12.5)] for $\Lambda_{L}$, as that could require power-saving estimates for linear correlations of the ideal counting function $a_{L}(n):=|\{\mathfrak{a}\text{ ideal of L}\ :\ \mathfrak{N}\mathfrak{a}=n\}|$. Such correlations could be studied using techniques of Browning and Matthiesen [7], but to achieve power-saving seems unlikely. Hence, we have decided to proceed in a different way.
@@ 2.6. Hooley’s method
Let $a\in\mathbb{Z}$ be not equal to $-1$ or a perfect square, and assume that $\operatorname{HRH}(a)$ holds true. Let $F:\mathbb{N}\to\mathbb{C}$ be a function with $|F(n)|\leq C$ for all $n\in\mathbb{N}$. Let $b\in\{0,\ldots,W-1\}$ and assume that $N\in\mathbb{N}$ is sufficiently large in terms of $a$. Then,
@@
Let $a\in\mathbb{Z}$ be not equal to $-1$ or a perfect square and assume $\operatorname{HRH}(a)$. Suppose that $w(N)$ grows to infinity with $N$ sufficiently slowly in terms of $a,m,d,Q,M$, and that $|\Delta_{a}|$ divides $\mathscr{D}$.
@@ 3.1. Elementary facts about maximal abelian subextensions
$[(\sigma_{1},\tau_{1}),(\sigma_{2},\tau_{2})]=([\sigma_{1},\sigma_{2}],[\tau_{1},\tau_{2}])=([\sigma_{1},\sigma_{2}],\operatorname{id})\in H,$
@@
$[(\sigma_{1},\tau_{1}),(\rho_{1},\lambda_{1})]\cdots[(\sigma_{k},\tau_{k}),(\rho_{k},\lambda_{k})]=([\sigma_{1},\rho_{1}]\cdots[\sigma_{k},\rho_{k}],[\tau_{1},\lambda_{1}]\cdots[\tau_{k},\lambda_{k}])=(\sigma,\operatorname{id}),$
@@ 3.2. Chebotarev classes
$[K:\mathbb{Q}][K(\mu_{q_{1}q_{2}}):\mathbb{Q}]=[K:\mathbb{Q}][K(\mu_{q_{1}}):\mathbb{Q}][\mathbb{Q}(\mu_{q_{2}}):\mathbb{Q}]=[K(\mu_{q_{1}}):\mathbb{Q}][K(\mu_{q_{2}}):\mathbb{Q}].$
@@
and thus $\tau_{\mathbf{K},\mathbf{C},\Psi}(p)=\beta_{\Psi,p}$. The remaining assertions now follow from [13, Lemma 1.3]. ∎
@@
$[K(\mu_{Q}):\mathbb{Q}(\mu_{Q})]=[K:K^{\operatorname{ab}}]=[K(\mu_{Qq}):\mathbb{Q}(\mu_{Qq})].$
@@ 3.3. Artin primes
$\beta_{a}(q)$ :=$:=\begin{cases}(-1)^{\frac{\frac{\Delta_{a}}{\gcd(q,\Delta_{a})}-1}{2}}\gcd(q,\Delta_{a}),&\mbox{if }\frac{\Delta_{a}}{\gcd(q,\Delta_{a})}\text{ is odd, }\\ 1&\mbox{otherwise, }\end{cases}$ (3.5)
@@
$\mathscr{A}_{a}(b,q):=\mathscr{A}_{a}\cdot\begin{cases}\frac{f_{a}^{\dagger}(q)}{\phi(q)}\prod_{p|b-1,p|q}\Big{(}1-\frac{1}{p}\Big{)},&\mbox{if }\gcd(b-1,q,h_{a})=\gcd(b,q)=1,\\ 0,&\mbox{otherwise. }\end{cases}$ (3.6)
@@
    Let $a\in\mathbb{Z}$ be not equal to $-1$ or a perfect square, and suppose that $|\Delta_{a}|\mid q_{1}$. Then
@@
Even though the execution is slightly more complicated, the basic idea of this proof is the same as in [13, Lemma 1.3]. We may assume $p$ to be sufficiently large so that $p\nmid h_{a_{1}}\cdots h_{a_{t}}\mathscr{D}_{\mathbf{a}}$ and such that no form is constant and no two of the linear parts of the forms $\psi_{i}$ are linearly dependent over $\mathbb{F}_{p}$. Recall from (1.7) that
@@
$S_{l}:=\{n\in(\mathbb{Z}/p\mathbb{Z})^{s}\ :\ \psi_{i}(n)\not\equiv 0\bmod p\ \text{ for all $i$ and }\ \psi_{i}(n)\equiv 1\bmod p\text{ for exactly }l\text{ indices }i\}.$
@@ 4. Equidistributed nilsequences
The following lemma can be thought of as a polynomial version of [15, Lemma 7.10]. However, even in the case $D=1$, our assumptions on the coefficients are slightly different. Recall the definition of the smoothness norms $\lVert p\rVert_{\operatorname{C^{\infty}}[N]}$ of polynomial sequences $p:\mathbb{Z}\to\mathbb{R}/\mathbb{Z}$ given in [15, Definition 2.7]: writing $p$ in the binomial basis as
@@
By our hypothesis, any $t=a_{0}^{i_{0}}\cdots a_{D}^{i_{D}}$ satisfies $|t|=|a_{0}^{i_{0}}\cdots a_{D}^{i_{D}}|\leq(QN)^{i_{0}+\cdots+i_{D}}N^{-i_{1}-2i_{2}-\cdots-Di_{D}}$. Hence, by Lemma 4.1, we obtain the estimate
@@
Next, we prove a simple weak converse to the quantitative Leibman theorem [15, Theorem 2.9]. For a similar statement, see also [29, Proposition 14.3].
@@
Let $m,d\geq 0$, then there is a constant $C(m,d)\geq 1$, such that the following holds. Let $20<\delta<1/$ 2 and $N\gg_{m,d}1$.
@@
This is inspired by arguments in the proofs of [15, Lemma A.17] and [14, Proposition 2.1]. By [15, Lemma 2.8], our assumption (4.8) implies that
@@
Let $x,y$ be as in the supremum above. Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [15, Lemma A.14], there are $x^{\prime}\in x$, $y^{\prime}\in y$ with $\psi(x^{\prime}),\psi(y^{\prime})\in[0,1)^{m}$. By the proof of [15, Lemma A.15], there is $\gamma\in\Gamma$, such that
@@
By [15, Lemma A.16], this quantity is $\ll_{m,d}\delta^{-O_{m,d}(1)}$. Using [15, Lemma A.4], we also have $d(x^{\prime},\operatorname{id}_{G})\ll_{m,d}\delta^{-O_{m,d}(1)}$, and by the triangle inequality $d(y^{\prime},\operatorname{id}_{G})\ll_{m,d}\delta^{-O_{m,d}(1)}$. Hence, [15, Lemma A.4] shows that
@@
The following proposition states, essentially, that polynomial subsequences of equidistributed polynomial nilsequences are again equidistributed polynomial nilsequences. For a similar result, see [29, Proposition 15.3].
@@
    The function $\tilde{g}:=g\circ P:\mathbb{Z}\to G$ is a polynomial sequence of degree $dD$. More precisely, $\tilde{g}\in\operatorname{poly}(\mathbb{Z},\tilde{G}_{\bullet})$, with the filtration $G=\tilde{G}_{0}=\tilde{G}_{1}\geq\cdots\geq\tilde{G}_{dD+1}=\{0\}$, where $\tilde{G}_{i}:=G_{\lceil i/D\rceil}$. Moreover, the Mal’cev basis $\mathscr{X}$ is also adapted to the filtration $\tilde{G}_{\bullet}$.
@@
    $|a_{i}|\leq\delta^{-c}N^{1-i/D}\quad\text{ for all }1\leq i\leq D,$
@@
For $i\in\mathbb{N}$, write $\tilde{m}_{i}:=\dim\tilde{G}_{i}=m_{\lceil i/D\rceil}$. With $H_{j}$ as in [15, Definition 2.1], we have $\tilde{G}_{i}=G_{\lceil i/D\rceil}=H_{m-\tilde{m}_{i}}$, hence $\mathscr{X}$ is also a Mal’cev basis adapted to the filtration $\tilde{G}_{\bullet}$.
Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [15, Lemma 6.7], we have
@@
For any $j\geq\lceil k/D\rceil$, we have $m_{j}\leq m_{\lceil k/D\rceil}=\tilde{m}_{k}$. Hence, if $i\leq m-\tilde{m}_{k}$, then also $i\leq m-m_{j}$ for all $j\geq\lceil k/D\rceil$, and thus $(\tilde{\mathbf{t}}_{k})_{i}=0$. Again by [15, Lemma 6.7], this shows that $p\circ P\in\operatorname{poly}(\mathbb{Z},\tilde{G}_{\bullet})$, as desired in (1).
@@
i.e. the sequence $(\bar{g}(n)\Gamma)_{n\in[\bar{N}]}$ with $\bar{g}(n)=\tilde{g}(u+vn)=g(P(u+vn))$ is not $\delta^{c}$-equidistributed. The quantitative Leibmann theorem [15, Theorem 2.9] yields a nontrivial horizontal character $\bar{\eta}:G\to\mathbb{R}/\mathbb{Z}$ with $0<|\bar{\eta}|\ll_{m,d}\delta^{-O_{m,d}(c)}$, such that
@@
Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [15, Lemma 6.7], we have
@@
as $k\leq m-m_{i}\leq m-m_{j}$ for all $j\geq i$. Again by [15, Lemma 6.7], this shows that $\tilde{g}\in\operatorname{poly}(\mathbb{Z},G_{\bullet})$, thus proving (1).
@@
In other words, the sequence $(\bar{g}(n)\Gamma)_{n\in[\bar{N}]}$ with $\bar{g}(n)=\tilde{g}(u+vn)$ is not $\delta^{c}$-equidistributed. The quantitative Leibmann theorem [15, Theorem 2.9] shows the existence of a nontrivial horizontal character $\bar{\eta}$ with $0<|\bar{\eta}|\ll_{m,d}\delta^{-O_{m,d}(c)}$, such that

### v3

@@ Acknowledgements
We thank Pierre-Yves Bienvenu for helpful discussions. C.F. was first exposed to the machinery invented by Green, Tao and Ziegler in a still ongoing collaboration with Lilian Matthiesen. In particular, he thanks Matthiesen for introducing him to her work on polynomial subsequences of equidistributed nilsequences. We thank Lilian Matthiesen and an anonymous referee for pointing out relevant references and bibliographic inaccuracies in an earlier version.
Most of this work was completed while M.T. was a visiting scholar at TU Graz in 2022–2024. The authors thank TU Graz for its hospitality. C.F. was supported by EPSRC grants EP/T01170X/1 and EP/T01170X/2. M.T. was supported by Czech Science Foundation (GAČR) grant 22-11563O.
@@ 2.1. Some notation
As already stated, for $N\in\mathbb{N}$ we let $N]:=\{1,\ldots,N\}$. We think of number fields as subfields of $\mathbb{C}$. We denote the discriminant of a number field $K$ by $\Delta_{K}$, the monoid of non-zero ideals of its ring of integers $\mathscr{O}_{K}$ by $\mathscr{I}_{K}$, the maximal abelian subextionsion of $K$ over $\mathbb{Q}$ by $K^{\operatorname{ab}}$, and the finite part of the conductor of this abelian subfield by $\Phi_{K^{\operatorname{ab}}}$. We denote the absolute norm of $\mathfrak{a}\in\mathscr{I}_{K}$ by $\mathfrak{N}\mathfrak{a}$.
@@ 2.3. Polynomial subsequences of equidistributed nilsequences
We prove our main results using the machinery developed by Green, Tao and Ziegler, in particular the inverse theorem for Gowers norms [17]. This makes it necessary to deal with polynomial sequences in nilmanifolds. Green and Tao have established the quantitative theory of polynomial nilsequences in [16]. We use the definitions and terminology of this work freely and will refer to it frequently. In particular, we use Mal’cev bases as defined in [16, Definition 2.1] and the metrics induced by these Mal’cev bases as defined in [16, Definition 2.2]. Lipschitz constants are understood with respect to this metric.
In Section 4, we compile some facts regarding polynomial subsequences of equidistributed nilsequences, which are probably well known to experts. Polynomial subsequences of nilsequences were first investigated by Matthiesen in [32, 33], and the main results in this section are technical variations of hers. We have included full proofs, as we were not able to find the exact required statements in the literature.
@@ 2.4. W-trick
A central aspect in the strategy of Green, Tao and Ziegler is the $W$-trick, which already appeared in [11, 12]. It requires us to consider an increasing function $w(N)$ that grows to infinity with $N$ and satisfies $1\leq w(N)\leq\frac{1}{3}\log\log N$, a constant $\mathscr{D}\in\mathbb{N}$, and the function
@@
$\Lambda^{\prime}(n)$ $:=\begin{cases}\log p,&\text{ if }n=p\text{ prime},\\ 0,&\text{ otherwise,}\end{cases}$
[…]
$\Lambda_{a}^{\prime}(n)$ $:=\begin{cases}\log p,&\text{ if }n=p\text{ prime with }\mathbb{F}_{p}^{\times}=\langle a\rangle,\\ 0,&\text{ otherwise.}\end{cases}$
@@
We will convince ourselves (Lemma 9.1) that the pseudorandom majorant constructed by Green and Tao in [14, Proposition 6.4] for $\Lambda_{b,W}$ is also sufficient for our $\Lambda_{K,C,b,W}$ and $\Lambda_{a,b,W}$. Standard arguments, including the generalised von Neumann Theorem [14, Proposition 7.1] (see Lemma 9.4), then reduce Theorem 1.4 to the following Gowers norm estimate. For the definition of $\lVert\cdot\rVert_{U^{u+1}[N]}$ we refer to [14, (B.13)].
@@
which was proved in [14, Theorem 7.2] with the choice of $\mathscr{D}=1$ and $w(N)=\log\log\log N$ but holds also in general (Lemma 9.2).
@@ 2.5. Von Mangoldt model and orthogonality with nilsequences
Green and Tao use the inverse theorem for Gowers norms [17] to deduce (2.3) from an orthogonality result of $\Lambda_{b,W}-1$ with nilsequences ([14, Proposition 10.2]). The proof of this result uses the identity
@@
Starting from this identity, Green and Tao reduce [14, Proposition 10.2] to a sieve estimate [14, (12.5)] and a strong orthogonality result for the Möbius function $\mu$ with nisequences, which they prove in [15, Theorem 1.1]. Unfortunately, an analog of (2.4) is not available for our functions $\Lambda_{K,C}$ and $\Lambda_{a}$.
Therefore, we proceed in a different way in our proofs of Proposition 2.1 and Proposition 2.2: in order to show orthogonality with nilsequences, we apply directly the strategy used for the Möbius function in [15]. This makes crucial use of the fact that $\mu(n)$ is equidistributed in arithmetic progressions, which is clearly not entirely true for our function $\Lambda_{K,C,b,W}(n)-\phi(W)\eta_{K,C}(b,W)$ due to progressions such as $n\equiv 0\bmod b$. To deal with this issue, we replace the constant average $\phi(W)\eta_{K,C}(b,W)$ by a suitably scaled version of the $W$-tricked classical von Mangoldt function $\Lambda_{b,W}(n)$ and show that this provides a good model for $\Lambda_{K,C,b,W}(n)$ in all arithmetic progressions. With this model, we show the following orthogonality result, which is analogous to [15, Theorem 1.1].
@@
$\ll_{K,m,d,A,\mathscr{D}}Q^{O_{m,d,A}(1)}(1+\lVert F\rVert_{\operatorname{Lip}})(\log N)^{-A}.$
@@
[…]
@@
We prove Theorem 2.4 in §5 and deduce Theorem 2.3 from it in §6. A generalisation of the results of Green and Tao to prime elements in number fields was obtained by Kai [23] and recently applied to Hilbert’s tenth problem by Koymans and Pagano [26]. Our Theorem 2.4 concerns norms of ideals and is independent of Kai’s work.
Let us just mention in passing that it would also be possible to deal with $\Lambda_{K,C}$ in terms of the von Mangoldt function $\Lambda_{L}$ for ideals in certain subfields $L$ of $K$ (see Lemma 6.2). For these ideal von Mangoldt functions, an analogue of (2.4) exists and our techniques from §5 would be capable of also proving an analogue of [15, Theorem 1.1] for the Möbius function $\mu_{L}$ of ideals of $L$. However, it seems quite difficult to prove the required analogue of the sieve estimate [14, (12.5)] for $\Lambda_{L}$, as that could require power-saving estimates for linear correlations of the ideal counting function $a_{L}(n):=|\{\mathfrak{a}\text{ ideal of L}\ :\ \mathfrak{N}\mathfrak{a}=n\}|$. Such correlations could be studied using techniques of Browning and Matthiesen [7], but to achieve power-saving seems unlikely. Hence, we have decided to proceed in a different way.
@@ 2.6. Hooley’s method
Let $a\in\mathbb{Z}$ be not equal to 1-1 or a perfect square, and assume that $\operatorname{HRH}(a)$ holds true. Let $F:\mathbb{N}\to\mathbb{C}$ be a function with $|F(n)|\leq C$ for all $n\in\mathbb{N}$. Let $b\in\{0,\ldots,W-1\}$ and assume that $N\in\mathbb{N}$ is sufficiently large in terms of $a$. Then,
@@
Let $a\in\mathbb{Z}$ be not equal to 1-1 or a perfect square and assume $\operatorname{HRH}(a)$. Suppose that $w(N)$ grows to infinity with $N$ sufficiently slowly in terms of $a,m,d,Q,M$, and that $|\Delta_{a}|$ divides $\mathscr{D}$.
@@ 3.1. Elementary facts about maximal abelian subextensions
$(\sigma_{1},\tau_{1}),(\sigma_{2},\tau_{2})]=([\sigma_{1},\sigma_{2}],[\tau_{1},\tau_{2}])=([\sigma_{1},\sigma_{2}],\operatorname{id})\in H,$
@@
$(\sigma_{1},\tau_{1}),(\rho_{1},\lambda_{1})]\cdots[(\sigma_{k},\tau_{k}),(\rho_{k},\lambda_{k})]=([\sigma_{1},\rho_{1}]\cdots[\sigma_{k},\rho_{k}],[\tau_{1},\lambda_{1}]\cdots[\tau_{k},\lambda_{k}])=(\sigma,\operatorname{id}),$
@@ 3.2. Chebotarev classes
$K:\mathbb{Q}][K(\mu_{q_{1}q_{2}}):\mathbb{Q}]=[K:\mathbb{Q}][K(\mu_{q_{1}}):\mathbb{Q}][\mathbb{Q}(\mu_{q_{2}}):\mathbb{Q}]=[K(\mu_{q_{1}}):\mathbb{Q}][K(\mu_{q_{2}}):\mathbb{Q}].$
@@
and thus $\tau_{\mathbf{K},\mathbf{C},\Psi}(p)=\beta_{\Psi,p}$. The remaining assertions now follow from [14, Lemma 1.3]. ∎
@@
$K(\mu_{Q}):\mathbb{Q}(\mu_{Q})]=[K:K^{\operatorname{ab}}]=[K(\mu_{Qq}):\mathbb{Q}(\mu_{Qq})].$
@@ 3.3. Artin primes
$\beta_{a}(q)$ $:=\begin{cases}(-1)^{\frac{\frac{\Delta_{a}}{\gcd(q,\Delta_{a})}-1}{2}}\gcd(q,\Delta_{a}),&\mbox{if }\frac{\Delta_{a}}{\gcd(q,\Delta_{a})}\text{ is odd, }\\ 1&\mbox{otherwise, }\end{cases}$ (3.5)
@@
[…]
@@
    Let $a\in\mathbb{Z}$ be not equal to 1-1 or a perfect square, and suppose that $|\Delta_{a}|\mid q_{1}$. Then
@@
Even though the execution is slightly more complicated, the basic idea of this proof is the same as in [14, Lemma 1.3]. We may assume $p$ to be sufficiently large so that $p\nmid h_{a_{1}}\cdots h_{a_{t}}\mathscr{D}_{\mathbf{a}}$ and such that no form is constant and no two of the linear parts of the forms $\psi_{i}$ are linearly dependent over $\mathbb{F}_{p}$. Recall from (1.7) that
@@
[…]
@@ 4. Equidistributed nilsequences
The following lemma can be thought of as a polynomial version of [16, Lemma 7.10]. However, even in the case $D=1$, our assumptions on the coefficients are slightly different. Recall the definition of the smoothness norms $\lVert p\rVert_{\operatorname{C^{\infty}}[N]}$ of polynomial sequences $p:\mathbb{Z}\to\mathbb{R}/\mathbb{Z}$ given in [16, Definition 2.7]: writing $p$ in the binomial basis as
@@
By our hypothesis, any $t=a_{0}^{i_{0}}\cdots a_{D}^{i_{D}}$ satisfies $|t|=|a_{0}^{i_{0}}\cdots a_{D}^{i_{D}}|\leq(QN)^{i_{0}+\cdots+i_{D}}N^{\frac{-i_{1}-2i_{2}-\cdots-Di_{D}}{D}}$. Hence, by Lemma 4.1, we obtain the estimate
@@
Next, we prove a simple weak converse to the quantitative Leibman theorem [16, Theorem 2.9]. That such a converse holds was first shown by Matthiesen in [32, Proposition 14.3], our statement here is a minor variation of this.
@@
Let $m,d\geq 0$, then there is a constant $C(m,d)\geq 1$, such that the following holds. Let $0<\delta<1/2$ and $N\gg_{m,d}1$.
@@
This is inspired by arguments in the proofs of [16, Lemma A.17] and [15, Proposition 2.1]. By [16, Lemma 2.8], our assumption (4.8) implies that
@@
Let $x,y$ be as in the supremum above. Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [16, Lemma A.14], there are $x^{\prime}\in x$, $y^{\prime}\in y$ with $\psi(x^{\prime}),\psi(y^{\prime})\in[0,1)^{m}$. By the proof of [16, Lemma A.15], there is $\gamma\in\Gamma$, such that
@@
By [16, Lemma A.16], this quantity is $\ll_{m,d}\delta^{-O_{m,d}(1)}$. Using [16, Lemma A.4], we also have $d(x^{\prime},\operatorname{id}_{G})\ll_{m,d}\delta^{-O_{m,d}(1)}$, and by the triangle inequality $d(y^{\prime},\operatorname{id}_{G})\ll_{m,d}\delta^{-O_{m,d}(1)}$. Hence, [16, Lemma A.4] shows that
@@
The following proposition states, essentially, that polynomial subsequences of equidistributed polynomial nilsequences are again equidistributed polynomial nilsequences. The first results concerning equidistribution of polynomial subsequences are due to Matthiesen [32, Proposition 15.3 and Proposition 15.4]. Our version is a minor technical variation of [32, Proposition 15.3], based on the same ideas.
@@
[…]
@@
    $|a_{i}|\leq\delta^{-c}N^{1-i/D}\quad\text{ for all }0\leq i\leq D,$
@@
For $i\in\mathbb{N}$, write $\tilde{m}_{i}:=\dim\tilde{G}_{i}=m_{\lceil i/D\rceil}$. With $H_{j}$ as in [16, Definition 2.1], we have $\tilde{G}_{i}=G_{\lceil i/D\rceil}=H_{m-\tilde{m}_{i}}$, hence $\mathscr{X}$ is also a Mal’cev basis adapted to the filtration $\tilde{G}_{\bullet}$.
Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [16, Lemma 6.7], we have
@@
[…]
@@
i.e. the sequence $(\bar{g}(n)\Gamma)_{n\in[\bar{N}]}$ with $\bar{g}(n)=\tilde{g}(u+vn)=g(P(u+vn))$ is not $\delta^{c}$-equidistributed. The quantitative Leibmann theorem [16, Theorem 2.9] yields a nontrivial horizontal character $\bar{\eta}:G\to\mathbb{R}/\mathbb{Z}$ with $0<|\bar{\eta}|\ll_{m,d}\delta^{-O_{m,d}(c)}$, such that
@@
Let $\psi:G\to\mathbb{R}^{m}$ denote the coordinates (of the second kind) with respect to our Mal’cev basis $\mathscr{X}$. By [16, Lemma 6.7], we have
@@
as $k\leq m-m_{i}\leq m-m_{j}$ for all $j\geq i$. Again by [16, Lemma 6.7], this shows that $\tilde{g}\in\operatorname{poly}(\mathbb{Z},G_{\bullet})$, thus proving (1).
@@
In other words, the sequence $(\bar{g}(n)\Gamma)_{n\in[\bar{N}]}$ with $\bar{g}(n)=\tilde{g}(u+vn)$ is not $\delta^{c}$-equidistributed. The quantitative Leibmann theorem [16, Theorem 2.9] shows the existence of a nontrivial horizontal character $\bar{\eta}$ with $0<|\bar{\eta}|\ll_{m,d}\delta^{-O_{m,d}(c)}$, such that